OpenStudy (brittany16):

Solve for the variable. 18 + y = 32

OpenStudy (anonymous):

subtract 18 from both sides

OpenStudy (anonymous):

y=14

jimthompson5910 (jim_thompson5910):

18 + y = 32 18 + y -18 = 32 - 18 y = 14